OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. (3 points) 4x2 − 12x + 9 16x2 + 24x − 9 4a2 − 10a + 25 36b2 − 24b − 16

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

I thought maybe the middle one.

OpenStudy (skullpatrol):

There is no middle one with 4 questions?

OpenStudy (anonymous):

oh, ahah sorry! I meant the 2nd one.

OpenStudy (anonymous):

the first one is perfect Square i.e (2x-3)^2

OpenStudy (skullpatrol):

$$ 4x^2 − 12x + 9$$ $$16x^2 + 24x − 9$$ $$4a^2 − 10a + 25$$ $$36b^2 − 24b − 16$$

OpenStudy (anonymous):

how did you find it though? I'd wrather learn than have the answer.

OpenStudy (skullpatrol):

@thisisarri GREAT answer @syedadnanshah

OpenStudy (skullpatrol):

Let's start with the definition of a perfect square trinomial

OpenStudy (anonymous):

ok, sure

OpenStudy (skullpatrol):

perfect square trinomial: An expression with three terms that is the square of a binomial.

OpenStudy (anonymous):

yeah. i know i just forgot how to find out which one it is(:

OpenStudy (skullpatrol):

What does the square of a binomial look like?

OpenStudy (anonymous):

by formula of (a-b)^2=a^2-2ab+b^2 your equation 4x^2-12x+9 can be wrtten as (2x)^2-2(2x)3+(3)^2 and so (2x-9)^2

OpenStudy (anonymous):

ok, but thats not one of the options? @syedadnanshah

OpenStudy (skullpatrol):

$$(a+b)^2= a^2 +2ab +b^2$$ or $$(a-b)^2=a^2-2ab+b^2$$

OpenStudy (anonymous):

that's the formula. @skullpatrol

OpenStudy (skullpatrol):

These are the forms you must work with.

OpenStudy (anonymous):

okay. So lets do this then(:

OpenStudy (anonymous):

whats's first @skullpatrol

OpenStudy (skullpatrol):

Lets look at $$4x^2-12x+9 $$

OpenStudy (anonymous):

good! ok

OpenStudy (anonymous):

i know hot so solve it

OpenStudy (anonymous):

*how

OpenStudy (skullpatrol):

You are not "solving" anything, you are simplifying an expression. In fact you are trying to factor it.

OpenStudy (anonymous):

yess, i know this. I'm just stuck on this one problem.

OpenStudy (skullpatrol):

Let me go to the drawing board...

OpenStudy (skullpatrol):